# Approximating the coefficients of the Bessel functions

### v3

@@ 9.3. Coefficients for the $D^{N}$ root system
If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
@@
See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10. Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$. ∎
@@
If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
@@ 10.1. Weak convergence to the free convolution
For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.15. Thus, part (E) of Lemma˜10.16 implies that for all $\lambda\in\Gamma$,
@@
the second equality follows from the fact that $J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}=\exp\\ \left(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+b_{\lambda}(N))p_{\lambda}\right)$ over $\Omega_{N}$ and $0\in\Omega_{N}$.
@@
[…]
@@ 10.2. Upper bounds of the magnitudes of Bessel functions
Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a,x\in\mathbb{C}^{N}$. Then,
$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}\leq\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}ha_{i})}.$
@@
$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}d\mu_{a}(\epsilon).$
@@
$e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}=e^{\sum_{i=1}^{N}\text{Re}(x_{i}\sum_{h\in H}c_{h}h^{\prime}ha_{i})}\leq\sum_{h\in H}c_{h}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}ha_{i})}$
@@
$\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}\epsilon_{i})}\leq\sum_{h\in H}\frac{c_{h}}{|H|}\sum_{h^{\prime}\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}h^{\prime}ha_{i})}=\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i}ha_{i})}.$
@@
Suppose $N\geq 1$ and $x,a\in\mathbb{C}^{N}$. Suppose $R$ and $I$ are nonnegative real numbers. Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$. Furthermore, assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
@@
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|\text{Re}(a_{i})|^{d}\leq R^{d}$ and $\frac{1}{N}\sum_{i=1}^{N}|\text{Im}(a_{i})|^{d}\leq I^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{R\sum_{i=1}^{N}|\text{Re}(x_{i})|+I\sum_{i=1}^{N}|\text{Im}(x_{i})|}.$
@@
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|\text{Re}(a_{i})|^{d}\leq d!R^{d}$ and $\frac{1}{N}\sum_{i=1}^{N}|\text{Im}(a_{i})|^{d}\leq d!I^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\left(\sum_{d=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d}\sum_{d=0}^{\infty}\left(2I\sum_{i=1}^{N}|\text{Im}(x_{i})|\right)^{d}\right)^{\frac{1}{2}}.$
@@
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}$ $=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\text{Re}(x),\text{Re}(\epsilon)\right\rangle-\left\langle\text{Im}(x),\text{Im}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)$
$\leq\left(\int_{\mathbb{C}^{N}}e^{2\left\langle\text{Re}(x),\text{Re}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)\int_{\mathbb{C}^{N}}e^{-2\left\langle\text{Im}(x),\text{Im}(\epsilon)\right\rangle}d\mu_{a}(\epsilon)\right)^{\frac{1}{2}}$
$\leq\left(\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\right)^{\frac{1}{2}}.$
@@
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{C}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(2|\text{Re}(x)|)M_{\nu}(|\text{Re}(\epsilon)|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
@@
$\frac{M_{\nu}(|\text{Re}(\epsilon)|)}{M_{\nu}(1,\ldots,1)}\leq\frac{\sum_{i=1}^{N}|\text{Re}(\epsilon_{i})|^{|\nu|}}{N}\leq\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|\text{Re}(ha_{i})|\right)^{|\nu|}}{N}.$
@@
$\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|\text{Re}(ha_{i})|\right)^{|\nu|}}{N}\leq\frac{1}{N}\sum_{h\in H}c_{h}\sum_{i=1}^{N}|\text{Re}(ha_{i})|^{|\nu|}=\frac{\sum_{i=1}^{N}|\text{Re}(a_{i})|^{|\nu|}}{N}\leq R^{|\nu|}.$
@@
(31) $\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(2|\text{Re}(x)|)}{\nu!}=e^{2R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
Similarly,
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq e^{2I\sum_{i=1}^{N}|\text{Im}(x)_{i}|}.$
Afterwards, it is straightforward to conclude the proof.
@@
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Re}(x)|,|\text{Re}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(2|\text{Re}(x)|)}{\pi(\nu)}=\sum_{k=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{k}.$
Similarly,
$\int_{\mathbb{C}^{N}}e^{2\left\langle|\text{Im}(x)|,|\text{Im}(\epsilon)|\right\rangle}d\mu_{a}(\epsilon)\leq\sum_{k=0}^{\infty}\left(2R\sum_{i=1}^{N}|\text{Im}(x_{i})|\right)^{k},$
@@
Lemma 10.6.
Suppose $N\geq 1$ and $x,a\in\mathbb{C}^{N}$. Suppose $M$ is a nonnegative real number. Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$. Furthermore, assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
  1. (A)
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a_{i}\right\rVert_{2}^{d}\leq M^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{M\sum_{i=1}^{N}\left\lVert x_{i}\right\rVert_{2}}.$
  2. (B)
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a_{i}\right\rVert_{2}^{d}\leq d!M^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\sum_{d=0}^{\infty}\left(M\sum_{i=1}^{N}\left\lVert x_{i}\right\rVert_{2}\right)^{d}.$
Proof.
For $z\in\mathbb{C}^{N}$, we let $\tilde{z}$ denote $(\left\lVert z_{i}\right\rVert_{2})_{i\in[N]}$. First, observe that for $a,b\in\mathbb{C}^{N}$,
$\left\lVert e^{\left\langle a,b\right\rangle}\right\rVert_{2}=e^{\text{Re}(\left\langle a,b\right\rangle)}\leq e^{\left\lVert\left\langle a,b\right\rangle\right\rVert_{2}}\leq e^{\left\langle\tilde{a},\tilde{b}\right\rangle}$.
Therefore, after referring to Theorem˜1.4, we can deduce that
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{C}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{C}^{N}}e^{\left\langle\tilde{x},\tilde{\epsilon}\right\rangle}d\mu_{a}(\epsilon).$
Afterwards, we can apply the same method used in the proof of Lemma˜10.5. ∎
@@
    Assume that $\theta\in\mathbb{C}$ such that $\mathcal{D}(D^{N}(\theta))$ is invertible. The value of $J_{a}^{D^{N}(\theta)}(z,0,\\ \ldots,0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}((N-1)\theta+\frac{1}{2})_{k}(\theta N)_{k}}z^{2k}$.
@@ 10.3. The uniform convergence of Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime
Suppose $m_{d}\in\mathbb{C}$ for $d\geq 1$. Assume that for $N\geq 1$, $a(N)\in\mathbb{C}^{N}$ such that $\lim_{N\rightarrow\infty}\frac{\sum_{i=1}^{N}a(N)_{i}^{d}}{N}=m_{d}$ for all $d\geq 1$. Suppose $M$ is a positive real number. Consider the following two conditions.
@@
    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a(N)_{i}\right\rVert_{2}^{d}\leq M^{d}$.
@@
    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}\left\lVert a(N)_{i}\right\rVert_{2}^{d}\leq d!M^{d}$.
@@
(A): By Lemma˜10.6, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
@@
Afterwards, by part (B) of Lemma˜10.8, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.

### v4

@@ 9.3. Coefficients for the $D^{N}$ root system
If $|\theta N|\rightarrow\infty$ and $\prod_{j=1}^{N}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, then
$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
@@
See (4), Lemma˜6.9, Lemma˜9.9, and Lemma˜9.10. Recall that for $\lambda,\nu\in\Gamma_{\text{even}}$, $\mathcal{M}^{D}_{\lambda\nu}\triangleq 2^{|\nu|-\ell(\nu)}\mathcal{M}^{A}_{\lambda\nu}$. Furthermore, we obtain the term $\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$ in $J_{a}^{D^{N}(\theta)}$ from $e,J_{a}^{D^{N}(\theta)}]_{D^{N}(\theta)}=e(a)$. ∎
@@
If $\theta N\rightarrow c\in\mathbb{C}$, $\prod_{j=1}^{N-k}(1+2(j-1)\theta)\not=0$ if $N$ is sufficiently large, and $2c$ is not a negative integer, then
$J^{D^{N}(\theta)}_{a}[0](x)=1+\frac{e(a)e(x)}{\prod_{j=1}^{N}(1+2(j-1)\theta)}$
and for $k\geq 1$ and $a,x\in\mathbb{C}^{N}$,
@@ 10.1. Weak convergence to the free convolution
For $N\geq 1$, $\mu_{a(N),\,b(N)}^{A^{N-1}(\theta)}$ is always compactly supported by Lemma˜10.1 and is therefore exponentially decaying, see Definition˜10.14. Thus, part (D) of Lemma˜10.15 implies that for all $\lambda\in\Gamma$,
@@
the second equality follows from the fact that $J_{a(N)}^{A^{N-1}(\theta)}J_{b(N)}^{A^{N-1}(\theta)}=\exp\big(\sum_{\lambda\in\Gamma_{N}}(a_{\lambda}(N)+\\ b_{\lambda}(N))p_{\lambda}\big)$ over $\Omega_{N}$ and $0\in\Omega_{N}$.
@@
the equation also holds after replacing $a_{(|B|)}$ with $b_{(|B|)}$. By the formula for the rectangular free convolution given in [rectangular_free_convolution]*Proposition 3.1, it is evident that $\Big\{\operatorname{\mathbb{E}}_{a\sim\mu_{a(N),\,b(N)}^{BC^{N}(\theta_{0},\theta_{1})}}\\ \left[\sum_{i=1}^{N}\frac{1}{2N}\delta\left(\frac{a_{i}}{\theta_{0}N}\right)+\frac{1}{2N}\delta\left(-\frac{a_{i}}{\theta_{0}N}\right)\right]\Big\}_{N\geq 1}$ converges in terms of moments to $\mu$. ∎
@@ 10.2. Upper bounds of the magnitudes of Bessel functions
Suppose $\theta\in\theta(\mathcal{R})$ is nonnegative. Suppose $a\in\mathbb{R}^{N}$ and $x\in\mathbb{C}^{N}$. Then,
$\left\lVert J^{\mathcal{R}(\theta)}_{a}(x)\right\rVert_{2}\leq\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i})ha_{i}}.$
@@
$\left\lVert J^{\mathcal{R}(\theta)}_{a(N)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}x_{i}h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}d\mu_{a}(\epsilon).$
@@
$e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}=e^{\sum_{i=1}^{N}\text{Re}(x_{i})\sum_{h\in H}c_{h}h^{\prime}ha_{i}}\leq\sum_{h\in H}c_{h}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}ha_{i}}$
@@
$\sum_{h^{\prime}\in H}\frac{1}{|H|}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}\epsilon_{i}}\leq\sum_{h\in H}\frac{c_{h}}{|H|}\sum_{h^{\prime}\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i})h^{\prime}ha_{i}}=\frac{1}{|H|}\sum_{h\in H}e^{\sum_{i=1}^{N}\text{Re}(x_{i})ha_{i}}.$
@@
Suppose $N\geq 1$, $a\in\mathbb{R}^{N}$, and $x\in\mathbb{C}^{N}$. Suppose $R$ is a nonnegative real number. Assume that $H(\mathcal{R})$ contains the reflections which permute the entries of $x$. Furthermore, assume that $\theta\in\theta(\mathcal{R})$ is nonnegative.
@@
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a_{i}|^{d}\leq R^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq e^{R\sum_{i=1}^{N}|\text{Re}(x_{i})|}.$
@@
    Assume that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a_{i}|^{d}\leq d!R^{d}$. Then,
    $\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}\leq\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d}.$
@@
$\left\lVert J_{a}^{\mathcal{R}(\theta)}(x)\right\rVert_{2}=\left\lVert\int_{\mathbb{R}^{N}}e^{\left\langle x,\epsilon\right\rangle}d\mu_{a}(\epsilon)\right\rVert_{2}\leq\int_{\mathbb{R}^{N}}e^{\left\langle\text{Re}(x),\epsilon\right\rangle}d\mu_{a}(\epsilon)\leq\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon).$
@@
$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)=\int_{\mathbb{R}^{N}}1+\sum_{\nu\in\Gamma}\frac{M_{\nu}(|\text{Re}(x)|)M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)\nu!}d\mu_{a}(\epsilon).$
@@
$\frac{M_{\nu}(|\epsilon|)}{M_{\nu}(1,\ldots,1)}\leq\frac{\sum_{i=1}^{N}|\epsilon_{i}|^{|\nu|}}{N}\leq\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|ha_{i}|\right)^{|\nu|}}{N}.$
@@
$\frac{\sum_{i=1}^{N}\left(\sum_{h\in H}c_{h}|ha_{i}|\right)^{|\nu|}}{N}\leq\frac{1}{N}\sum_{h\in H}c_{h}\sum_{i=1}^{N}|ha_{i}|^{|\nu|}=\frac{\sum_{i=1}^{N}|a_{i}|^{|\nu|}}{N}\leq R^{|\nu|}.$
@@
(31) $\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\nu!}=e^{R\sum_{i=1}^{N}|\text{Re}(x)_{i}|}.$
@@
$\int_{\mathbb{R}^{N}}e^{\left\langle|\text{Re}(x)|,|\epsilon|\right\rangle}d\mu_{a}(\epsilon)\leq 1+\sum_{\nu\in\Gamma}|\nu|!R^{|\nu|}\frac{M_{\nu}(|\text{Re}(x)|)}{\pi(\nu)}=\sum_{d=0}^{\infty}\left(R\sum_{i=1}^{N}|\text{Re}(x_{i})|\right)^{d},$
@@
    Assume that $\theta\in\mathbb{C}$ such that $\mathcal{D}(D^{N}(\theta))$ is invertible. The value of $J_{a}^{D^{N}(\theta)}(z,0,\ldots,\\ 0)$ is $1+\sum_{k=1}^{\infty}\frac{c_{k}(a^{2})}{4^{k}((N-1)\theta+\frac{1}{2})_{k}(\theta N)_{k}}z^{2k}$.
@@ 10.3. The uniform convergence of Bessel functions in the $\theta N\rightarrow c\in\mathbb{C}$ regime
Suppose $m_{d}\in\mathbb{R}$ for $d\geq 1$. Assume that for $N\geq 1$, $a(N)\in\mathbb{R}^{N}$ such that $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}a(N)_{i}^{d}=m_{d}$ for all $d\geq 1$. Suppose $M$ is a positive real number. Consider the following two conditions.
@@
    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a(N)_{i}|^{d}\leq M^{d}$.
@@
    For sufficiently large $N\in\mathbb{N}$, it is the case that for all $d\in\mathbb{N}$, $\frac{1}{N}\sum_{i=1}^{N}|a(N)_{i}|^{d}\leq d!M^{d}$.
@@
(A): By Lemma˜10.5, $\zeta\triangleq\{J^{A^{N-1}(\theta)}_{a(N)}(z_{1},\ldots,z_{r},0,\ldots,0)\}_{N\geq r}$ is uniformly bounded over compact subsets of $\Omega$, where $\Omega=\mathbb{C}^{r}$ if (C1) is satisfied and $\Omega=\Omega_{r}$ if (C2) is satisfied.
@@
Afterwards, by part (B) of Lemma˜10.7, we have that $\zeta$ converges to $\exp\left(\sum_{k\geq 1}c_{(k)}p_{(k)}\right)$ uniformly over compact subsets of $\Omega$.